OpenStudy (anonymous):

Simplify cube root of 5 over fourth root of 5? Medals 0 5 to the power of 1 over 4 5 to the power of 1 over 12 5 to the power of 7 over 12 5 to the power of 4 over 3 @DanJS

OpenStudy (danjs):

k

OpenStudy (danjs):

let me type it out

OpenStudy (danjs):

\[\frac{ 5^{1/3} }{ 5^{1/4} } = 5^{1/3 - 1/4}\]

OpenStudy (danjs):

1/3 - 1/4 = 1/12

OpenStudy (danjs):

\[5^{1/12}\]

OpenStudy (danjs):

When you move a power to the other side of a fraction, you change the sign on the power. When you multiply two like bases raised to powers, you add the powers.

OpenStudy (danjs):

THose are the two rules

OpenStudy (danjs):

cubed root means - to the 1/3 power fourth root means- to the 1/4 power nth root means - to the 1/n power

OpenStudy (anonymous):

ok

OpenStudy (danjs):

any others?

OpenStudy (anonymous):

Not yet.. so what answer choice would that be?

OpenStudy (danjs):

5^(1/12) 5 to the power of 1 over 12